# Defects in Wigner crystals: fracton-elasticity duality and vacancy proliferation

### v1

@@ I Introduction
The program outlined above is carried out using the tools of fracton-elasticity duality. Fractons – emergent particles with nontrivial constraints on their mobility [9, 10, 11, 12] – occupy a special place among the various types of quasiparticle excitations with which condensed matter physics abounds. One of the many exotic features of fracton systems is the emergence of unusal “higher-rank” gauge fields [13, 14, 15, 16], which mediate interactions between the fracton particles. While fractons have been theorized to emerge in a diverse range of platforms, one of their best-known realizations is as elastic defects in two-dimensional crystals: there, defects called disclinations are known to be immobile and disclination dipoles (i.e., dislocations) able to move along the dipole direction only. Simultaneously, the aforementioned higher-rank gauge fields have been shown to arise as duals of the elastic degrees of freedom [17, 18, 19, 20, 21, 13, 22, 23, 24].
Following the pioneering work on two-dimensional crystals, different versions of the fracton-elasticity duality have also been proposed for three-dimensional crystals [25, 26], Cosserat (also known as micropolar) crystals [27, 28], quasicrystals [29, 30], supersolids [31, 32], and vortex lattices [33, 34, 35]. The present work extends this list by deriving the duality for crystals composed of electrically charged particles. Although some generalisations of the fracton-elasticity duality to charged crystals are present in the literature [20, 25], these works suffer a serious limitation: while they take into account some defects, namely dislocations and disclinations, they entirely neglect vacancies and interstitials. The reason for this omission is that while disclinations and dislocations are topological defects, vacancies/interstitials in regular (uncharged) crystals are not, making them difficult to define and treat analytically [36, 37]. At the same time, however, these defects might be the most important ones to take into account. Indeed, while the energy cost of a dislocation or a disclination grows with the system size, the energy needed to create a vacancy or an interstitial is finite [38, 39, 40], meaning that a finite density of them will be present at any non-zero temperature. Moreover, some works [41, 42, 43] suggest that even in zero temperature, the two-dimensional Wigner crystal can undergo self-doping such that a finite density of interstitials is present in the ground state.
@@ III Effective action
    where $\mathbf{x}_{j}(t)$ is the position of site $j$ at time $t$ and $N$ is the total number of sites. We then split the full gauge field $\kappa$ into $\kappa^{\mathrm{latt}}$ and the remaining part $\phi(x)$ [^3] An similar splitting was performed for a vortex lattice in superfluids in Ref. [54].,
@@
The building blocks of the effective action are the combinations of $u_{i}$ and $\phi$ that are invariant under $T_{i}$ and $\mathcal{Q}^{(1)}$. Such combinations are
@@ IV Dual gauge theory
Crystal defects give rise to singularities in the low-energy fields $u_{i}$ and $\phi$. Thus, in order to study their behavior more closely, we decompose the fields into smooth parts denoted with a bar and singular parts denoted with the superscript $s$:
@@
Since the full gauge field $\kappa$ is the dual electromagnetic $U(1)$ gauge field, it exhibits point singularities at the locations of electric monopoles. At the same time, given the splitting $\kappa=\kappa^{\mathrm{latt}}+\phi$ (see Sec. III), the singularities associated with the lattice sites and the background charge are incorporated into $\kappa^{\mathrm{latt}}$. Consequently, $\phi$ is regular everywhere except at the lattice sites at which an electron is missing, such as the vacancy site in Fig. 1, or when an excess electron forms an interstitial. To be more precise, if we define the vacancy density as (in the vector calculus notation)
$\rho^{\mathrm{vac}}\equiv\nabla\cdot\left(\nabla\times\boldsymbol{\phi}^{s}\right),$ (17)
@@
$\partial_{\nu}J^{\mathrm{disc}}_{i\mu\nu}$ $=0,$ (34a)
$\partial_{\nu}J^{\mathrm{disl}}_{i\mu\nu}$ $=0,$ (34b)
$\partial_{t}\left(\rho^{\mathrm{vac}}-n_{0}\epsilon_{ijk}x_{i}\rho^{\mathrm{disl}}_{jk}\right)-\partial_{l}\left(n_{0}\epsilon_{ijk}x_{i}J^{\mathrm{disl}}_{jkl}\right)$ $=0.$ (34c)
@@
  • Integrating the $\mu=t$ component of Eq. (34a) ((34b)) over a finite volume, we find that the total Frank vector (Burgers vector) of defects crossing any closed compact surface is zero. Thus, disclinations (dislocations) must form closed loops or extend to the boundaries of the system.
  • Calculating the surface integral of Eq. (34a) ((34b)) over a closed spatial surface, we find that the total Frank vector (Burgers vector) of defects crossing any closed surface is conserved. For example, assuming that the auxiliary surface in Fig. 3 extends to infinity, its total Frank vector and Burgers vector remain constant in time.
  • Eq. (34c) expresses the conservation of the total electric charge. It implies that a dislocation line is free to move in the plane spanned by the Burgers vector and the direction of the dislocation line $\boldsymbol{\rho}^{\mathrm{disl}}_{i}$, motion known in the theory of elasticity as a glide. On the other hand, motion perpendicular to that plane, known as a climb, has to be accompanied by absorbing or emitting a vacancy.
@@ V Interactions of vacancies and interstitials
$E_{\mathrm{int}}=\frac{C_{s}}{2n_{0}^{2}}\frac{Q_{1}Q_{2}}{R^{3}}.$ (43)
@@ VI Condensation of vacancies or interstitials
Mode dispersions can be found after deriving the equations of motion from the action (47). In App. (S3), we pursue another approach, in which the remaining gapped modes are first integrated out, revealing more directly the low-energy structure of the effective action. Either way, the massless spectrum is found to be as follows.
@@
These results suggest the following behavior of the longitudinal mode. In the Coulomb phase, longitudinal oscillations have a gap equal to the plasma frequency, as displacing the lattice electrons (as in Fig. 2) results in charge separation, leading to finite-frequency oscillations. In this phase, lattice compressions are not allowed at low frequencies. On the other hand, in the presence of a finite density of point defects, one can imagine periodic lattice compressions accompanied by a redistribution of the vacancies/interstitials in such a way that the coarse-grained charged density remains zero everywhere. This way, plasma oscillations are avoided and long-wavelength longitudinal waves can be present. Based on this intuition, we speculate that the appearance of a longitudinal mode is a generic scenario for charged crystals in the presence of a finite density of charged defects. At the same time, in the extreme case of a fully melted crystal, all the modes are gapped by plasma oscillations [63], meaning that a coexistence of the elastic lattice and the charged defects is crucial for the presence of gapless sound waves.
@@ VIII Acknowledgements
I thank Tobias Holder, Francisco Peña-Benítez, Aleksander Głódkowski and Lazaros Tzaloukidis for comments on the manuscript. I acknowledge financial support by the European Research Council (ERC) under grant QuantumCUSP (Grant Agreement No. 101077020) and by the Deutsche Forschungsgemein- schaft (DFG) through the cluster of excellence ct.qmat (EXC 2147, project-id 390858490).
@@
The continuity equations (34a) and (34b) are the different components of Eq. (S19), while Eq. (34c) can be obtained by combining (S20) with (34b).
@@
Varying the action (S26) with respect to $A^{(\Theta)}_{ijk}$ gives combinations of derivatives of the Ampère’s law, and therefore provides no new information. To proceed, we now use the gauge freedom to set $A^{(\Theta)}_{i}=0$, such that $F_{ij}$ and $\tau_{\langle ij\rangle}$ are given in terms of derivatives of $\tilde{A}_{\langle ij\rangle}$ only. Furthermore, assuming that the perturbation propagates along the $x$ direction, we get from Eq. (S29) $\tilde{A}_{\langle yy\rangle}=\tilde{A}_{\langle zz\rangle}=\tilde{A}_{\langle yz\rangle}=0$, leaving $\tilde{A}_{\langle xy\rangle}$ and $\tilde{A}_{\langle xz\rangle}$ as the only propagating degrees of freedom. The dispersion relation of both modes reads

### v2

@@ I Introduction
The program outlined above is carried out using the tools of fracton-elasticity duality. Fractons – emergent particles with nontrivial constraints on their mobility [40, 49, 25, 22] – occupy a special place among the various types of quasiparticle excitations with which condensed matter physics abounds. One of the many exotic features of fracton systems is the emergence of unusal “higher-rank” gauge fields [53, 24, 47, 30], which mediate interactions between the fracton particles. While fractons have been theorized to emerge in a diverse range of platforms, one of their best-known realizations is as elastic defects in two-dimensional crystals: there, defects called disclinations are known to be immobile and disclination dipoles (i.e., dislocations) able to move along the dipole direction only. Simultaneously, the aforementioned higher-rank gauge fields have been shown to arise as duals of the elastic degrees of freedom [32, 33, 72, 3, 50, 53, 52, 6, 65, 66].
Following the pioneering work on two-dimensional crystals, different versions of the fracton-elasticity duality have also been proposed for three-dimensional crystals [4, 45], Cosserat (also known as micropolar) crystals [23, 26], quasicrystals [60, 17], supersolids [51, 35], and vortex lattices [41, 42, 13]. The present work extends this list by deriving the duality for crystals composed of electrically charged particles. Although some generalisations of the fracton-elasticity duality to charged crystals are present in the literature [3, 4], these works suffer a serious limitation: while they take into account some defects, namely dislocations and disclinations, they entirely neglect vacancies and interstitials. The reason for this omission is that while disclinations and dislocations are topological defects, vacancies/interstitials in regular (uncharged) crystals are not, making them difficult to define and treat analytically [39, 11]. At the same time, however, these defects might be the most important ones to take into account. Indeed, while the energy cost of a dislocation or a disclination grows with the system size, the energy needed to create a vacancy or an interstitial is finite [15, 10, 7], meaning that a finite density of them will be present at any non-zero temperature. Moreover, some works [2, 46, 31] suggest that even at zero temperature, the two-dimensional Wigner crystal can undergo self-doping such that a finite density of interstitials is present in the ground state.
@@ III Effective action
    where $\mathbf{x}_{j}(t)$ is the position of site $j$ at time $t$ and $N$ is the total number of sites. We then split the full gauge field $\kappa$ into $\kappa^{\mathrm{latt}}$ and the remaining part $\phi$, which just like $\kappa$ is a one-form with one temporal and three spatial components $\phi=\phi_{t}dt+\phi_{i}dx_{i}$ [^3] A similar splitting was performed for a vortex lattice in superfluids in Ref. [69].:
@@
The building blocks of the effective action are the combinations of $u_{i}$ and $\phi_{\mu}$ that are invariant under $T_{i}$ and $\mathcal{Q}^{(1)}$. Such combinations are
@@ IV Dual gauge theory
Crystal defects give rise to singularities in the low-energy fields $u_{i}$ and $\phi=\phi_{t}dt+\phi_{i}dx_{i}$. Thus, in order to study their behavior more closely, we decompose the fields into smooth parts denoted with a bar and singular parts denoted with the superscript $s$:
@@
Since the full gauge field $\kappa$ is the dual electromagnetic $U(1)$ gauge field, it exhibits point singularities at the locations of electric monopoles. At the same time, given the splitting $\kappa=\kappa^{\mathrm{latt}}+\phi$ (see Sec. III), the singularities associated with the lattice sites and the background charge are incorporated into $\kappa^{\mathrm{latt}}$. Consequently, $\phi$ is regular everywhere except at the lattice sites at which an electron is missing, such as the vacancy site in Fig. 1, or when an excess electron forms an interstitial. To be more precise, if we define the vacancy density as
$\rho^{\mathrm{vac}}\equiv\partial_{i}\left(\epsilon_{ijk}\partial_{j}\phi^{s}_{k}\right),$ (17)
@@
$\begin{split}&\partial_{t}\left(\rho^{\mathrm{vac}}-n_{0}\epsilon_{ijk}x_{i}\rho^{\mathrm{disl}}_{jk}\right)\qquad\qquad\qquad\\ &\qquad\qquad+\partial_{i}\left(J^{\mathrm{vac}}_{i}-n_{0}\epsilon_{jkl}x_{j}J^{\mathrm{disl}}_{kli}\right)=0,\end{split}$ (34a)
$\partial_{\nu}J^{\mathrm{disc}}_{i\mu\nu}=0,$ (34b)
$\partial_{\nu}J^{\mathrm{disl}}_{i\mu\nu}=0.$ (34c)
@@
  • Integrating the $\mu=t$ component of Eq. (34b) ((34c)) over a finite volume, we find that the total Frank vector (Burgers vector) of defects crossing any closed compact surface is zero. Thus, disclinations (dislocations) must form closed loops or extend to the boundaries of the system.
  • Calculating the surface integral of Eq. (34b) ((34c)) over a closed spatial surface, we find that the total Frank vector (Burgers vector) of defects crossing any closed surface is conserved. For example, assuming that the auxiliary surface in Fig. 3 extends to infinity, its total Frank vector and Burgers vector remain constant in time.
  • Eq. (34a) expresses the conservation of the total electric charge. It implies that a dislocation line is free to move in the plane spanned by the Burgers vector and the direction of the dislocation line $\boldsymbol{\rho}^{\mathrm{disl}}_{i}$, motion known in the theory of elasticity as a glide. On the other hand, motion perpendicular to that plane, known as a climb, has to be accompanied by absorbing or emitting a vacancy.
@@ V Interactions of vacancies and interstitials
$E_{\mathrm{int}}=\frac{C_{s}}{2\pi n_{0}^{2}}\frac{Q_{1}Q_{2}}{R^{3}}.$ (43)
@@ VI Condensation of vacancies or interstitials
Mode dispersions can be found after deriving the equations of motion from the action (47). In App. S3, we pursue another approach, in which the remaining gapped modes are first integrated out, revealing more directly the low-energy structure of the effective action. Either way, the massless spectrum is found to be as follows.
@@
These results suggest the following behavior of the longitudinal mode. In the Coulomb phase, longitudinal oscillations have a gap equal to the plasma frequency, as displacing the lattice electrons (as in Fig. 2) results in charge separation, leading to finite-frequency oscillations. In this phase, lattice compressions are not allowed at low frequencies. On the other hand, in the presence of a finite density of point defects, one can imagine periodic lattice compressions accompanied by a redistribution of the vacancies/interstitials in such a way that the coarse-grained charged density remains zero everywhere. This way, plasma oscillations are avoided and long-wavelength longitudinal waves can be present. Based on this intuition, we speculate that the appearance of a longitudinal mode is a generic scenario for charged crystals in the presence of a finite density of mobile charged defects. At the same time, in the extreme case of a fully melted crystal, all the modes are gapped by plasma oscillations [67], meaning that a coexistence of the elastic lattice and the charged defects is crucial for the presence of gapless sound waves.
@@
The scenario discussed here is qualitatively different from the vacancy condensation in superfluid vortex crystals analyzed in [41]. In the latter case, the breaking of time-reversal symmetry by the background magnetic field introduces a gyroscopic term proportional to $\epsilon_{ij}u_{i}\partial_{t}u_{j}$ in the (2+1)-dimensional Lagrangian. This term, which has no counterpart in our action (11), renders $u_{x}$ and $u_{y}$ as a pair of canonically conjugate variables. Consequently, the vortex crystal supports only a single gapless mode with a quadratic dispersion, preventing the anti-Higgs mechanism from occurring.
@@ VIII Acknowledgements
I thank Tobias Holder, Francisco Peña-Benítez, Aleksander Głódkowski and Lazaros Tsaloukidis for comments on the manuscript. I acknowledge financial support by the European Research Council (ERC) under grant QuantumCUSP (Grant Agreement No. 101077020) and by the Deutsche Forschungsgemein- schaft (DFG) through the cluster of excellence ct.qmat (EXC 2147, project-id 390858490).
@@
The continuity equations (34b) and (34c) are the different components of Eq. (S19), while Eq. (34a) can be obtained by combining (S20) with (34c).
@@
Varying the action (S26) with respect to $A^{(\Theta)}_{ijt}$ gives the “Gauss’s law”
@@
Varying the action (S26) with respect to $A^{(\Theta)}_{ijk}$ gives combinations of derivatives of the Ampère’s law, and therefore provides no new information. Incidentally, this shows that the low-energy gauge theory can be simplified by gauge-fixing $A^{(\Theta)}_{ijk}=0$ before calculating variations of the action. Similarly, since Eq. (S29) is symmetric in $i$ and $j$, gauge-fixing $A^{(\Theta)}_{[ij]t}=0$ at the level of the action does not change the equations of motion either. Thus, an equivalent but simpler gauge theory can be formulated using only the traceless symmetric tensor $\tilde{A}_{\langle ij\rangle}$ and the symmetric tensor $A^{(\Theta)}_{(ij)t}$ with the definitions
[…]
The gauge transformations of the remaining gauge fields are parametrized by the symmetric tensor $\lambda^{(\Theta)}_{(ij)}$ and the vector $\lambda^{(\Theta)}_{i0}$ and read
[…]
After finding the equations of motion (S28) and (S29), we can further gauge-fix $A^{(\Theta)}_{ijt}=0$, such that $F_{ij}$ and $\tau_{\langle ij\rangle}$ are given in terms of derivatives of $\tilde{A}_{\langle ij\rangle}$ only. Furthermore, assuming that the perturbation propagates along the $x$ direction, we get from Eq. (S29) $\tilde{A}_{\langle yy\rangle}=\tilde{A}_{\langle zz\rangle}=\tilde{A}_{\langle yz\rangle}=0$, leaving $\tilde{A}_{\langle xy\rangle}$ and $\tilde{A}_{\langle xz\rangle}$ as the only propagating degrees of freedom. The dispersion relation of both modes reads